Please help me
1. find the quotient and write the the answer
in scientific notation (10*10^6)/5x10^-3=
2. Add 46x^2-100x-120 to 26x^2-49x+70=
3. write this number in scientific notation 41,700= ___ *10^.
Thank you so much.

Answer by lwsshak3(11628)

1. find the quotient and write the the answer
in scientific notation (10*10^6)/5x10^-3=2*10^9
2. Add 46x^2-100x-120 to 26x^2-49x+70=72x^2-149x-50
3. write this number in scientific notation 41,700= 417 *10^2